# Constructing Concise Characteristic Samples for Acceptors of Omega Regular Languages

### v3

@@ 3.3. Polynomial Time Learners and Teachers
A class $\mathbb{C}$ has concise characteristic samples if there exist a polynomial $p(n)$, a teacher $\mathbf{T}$, and a learner $\mathbf{L}$ such that for every $\mathcal{C}\in\mathbb{C}$, $\mathbf{T}(\mathcal{C})$ is a characteristic sample for $\mathcal{C}$ and $\mathbf{L}$, and $|\mathbf{T}(\mathcal{C})|\leq p(\textit{size}(\mathcal{C}))$. Recall that the length of a sample is the sum of length of the strings in it. In Corollary 3.6 we considered just the cardinality of $\mathbf{T}(\mathcal{C})$ and not the length of the words in the sample, but in $|\mathbf{T}(\mathcal{C})|$ the length of the words in the sample matter. Note also that the polynomial bound is in terms of the size of the smallest representation of ${\llbracket}\mathcal{C}{\rrbracket}$. If $\mathbb{C}$ has concise characteristic samples then $\mathbb{C}$ is concisely distinguishable, by Lemma 3.1. In the converse direction, we have the following consequence of Corollary 3.6.
@@
Assume the class $\mathbb{C}$ satisfies (C1), (C2), and (C3) and is concisely distinguishable. Then it has concise characteristic samples for a computable teacher $\mathbf{T}$ and a computable learner $\mathbf{L}$.
@@
As has been observed before [Pit89], if (P1) and (P2) are satisfied, this definition is superfluous because the learning algorithm of identification by enumeration used in the proof of Theorem 3.2 can be modified to run in polynomial time and the teacher adjusted appropriately to create a (potentially ridiculously large) characteristic sample. This observation motivates the introduction of the criterion of identification in the limit with polynomial time and data (described in Sec. 3.4). The argument for this observation follows.
@@ 3.5. Relation to learning with equivalence and membership queries
Suppose the class $\mathbb{C}$ satisfies properties (P1), (P2), and (P3). If $\mathbb{C}$ is polynomially learnable using membership and equivalence queries, then it is also efficiently teachable.
@@ 4.2. Negative results for nondeterministic classes
The classes $\mathbb{NBA}$, $\mathbb{NPA}$, $\mathbb{NMA}$, $\mathbb{NCA}$ do not have concise characteristic sets [AFS20]. The proof is by constructing a family of languages $\{L_{n}\}_{n\in\mathbb{N}}$ with an acceptor of size quadratic in $n$ for which at least one word of length at least exponential in $n$ must be included in any characteristic sample for $L_{n}$. [^7] A negative result regarding query learning of $\mathbb{NBA}$, $\mathbb{NPA}$, and $\mathbb{NMA}$ was obtained by Angluin et al. [AAF20]. That result makes a plausible assumption of cryptographic hardness, which is not required here. Since an NBA (resp. NCA) is a special case of NRA (resp. NSA) the same is true for $\mathbb{NRA}$ and $\mathbb{NSA}$.
@@ 4.3. Consequences of membership and equivalence algorithms
Let $\mathbf{T}$ be a teacher witnessing the efficient teachability of M2MAs. The teacher $\mathbf{T}^{\prime}$ with input a SUBA $\mathcal{C}$ generates the characteristic sample $\mathbf{T}(r(\mathcal{C}))$, to which it adds one example $(x,b)$ such that $x=\mathcal{C}$ and $b=1$ iff $x\in{\llbracket}\mathcal{C}{\rrbracket}$. The learner $\mathbf{L}$ with input $T$ searches for the least $x$ (if any) such that $(x,b)\in T$, $r(x)$ is consistent with $T$, and $\mathbf{T}(r(x))\subseteq T$. If such an $x$ is found, it is output; otherwise, $\mathbf{L}$ constructs and outputs a SUBA consistent with $T$. A procedure constructing a default SUBA acceptor that agrees with a given sample $T$ is given in the proof of Proposition 5.2.
@@ 5.1. Duality
There are reductions of the problem of efficient teachability between $\mathbb{IBA}$ and $\mathbb{ICA}$ and between $\mathbb{IRA}$ and $\mathbb{ISA}$, using the duality between these types of acceptors. Consequently we focus on the classes $\mathbb{IBA}$, $\mathbb{IPA}$, $\mathbb{IRA}$, and $\mathbb{IMA}$ in what follows. [^8] The results regarding the classes $\mathbb{IBA}$ (and $\mathbb{ICA}$), $\mathbb{IPA}$, and $\mathbb{IMA}$ were obtained in [AFS20]; here we extend them to the classes $\mathbb{IRA}$ (and $\mathbb{ISA}$). Results for identifiability in the limit using polynomial time and data (but not efficient teachability) of the classes $\mathbb{IRA}$ (and $\mathbb{ISA}$) have also been provided in [BL21] using a different algorithm.
@@
Let $\mathcal{A}$ be an ICA. Because $\mathcal{A}$ is deterministic and complete, if we let $\mathcal{A}^{\prime}$ denote the IBA with the same components as $\mathcal{A}$, then $\mathcal{A}^{\prime}$ accepts the complement of the language $\mathcal{A}$, by Claim 4.
@@
The same conversion may be done with acceptors of types IRA and ISA, by Claim 5. ∎
@@ 5.2. The default acceptor
Given a sample $T$, we initialize $U$ to be the empty set, and for every word $w_{i}$ in the sample, we find the shortest prefix $u_{i}$ of $w_{i}$ that distinguishes it from all other examples in $T$ and add it to $U$ (see Fig. 2, middle). We arrange the finite words in $U$ in a trie in the usual manner. We add self-loops on each $\sigma\in\Sigma$ to the leaves of the trie (see Fig. 2, left). This deterministic automaton is termed the prefix-tree automaton [OG92]. If the automaton is incomplete, we add a new dead state with self-transitions on each $\sigma\in\Sigma$, and define all undefined transitions to go to the dead state. Recall that by Proposition 4.1 the length of a prefix distinguishing two examples $u_{1}(v_{1})^{\omega}$ and $u_{2}(v_{2})^{\omega}$ is polynomially bounded by the length of the examples. It follows that the prefix-tree automaton can be constructed in time polynomial in the length of the sample $T$.
@@
For a DBA, the acceptance condition $F$ consists of all the trie-leaf states that are prefixes of positive examples in $T$ (see Fig. 2, left). For a DMA, the acceptance condition consists of $\{\{q\}\mid q\in F\}$. For a DCA, the acceptance condition consists of the dead state (if one was added) and all the trie-leaf states that are not prefixes of positive examples in $T$. The DBA thus constructed may be transformed to a DPA or a DRA using Claim 1 or Claim 2, and the DCA may be transformed to a DSA using Claim 3.
@@
For a SUBA the construction is different. We first construct a set $V$ of shortest periods of positively labeled examples, by going iteratively over the examples, and for example $w_{i}=u(v)^{\omega}$ if $v_{i}$ is a shortest period of $w_{i}$, and none of the rotations of $v_{i}$ is in $V$ we add $v_{i}$ to $V$. Then for every period $v=\sigma_{1}\sigma_{2}\ldots\sigma_{k}$ in $V$ we construct an automaton with $k$ states $s_{0},s_{1},\ldots,s_{k-1}$ and transitions $(s_{i},\sigma_{i+1},s_{i^{\prime}})$ for $0\leq i<k$ where $i^{\prime}=(i+1)\mod k$. We designate $s_{0}$ as an accepting state. Finally, for each positively labeled word $w$ we look for the shortest prefix $x$ of $w$ such that $w=x(y)^{\omega}$ for some rotation $y=\sigma_{i}\ldots\sigma_{k}\sigma_{1}\ldots\sigma_{i-1}$ of a period $v=\sigma_{1}\sigma_{2}\ldots\sigma_{k}\in V$. For a rotation $y$ of $v\in V$, let $X_{y}$ be the set of such shortest prefixes. The prefixes in $X_{y}$ are arranged in a suffix-sharing trie and the trie is connected to the automata constructed for the periods by landing in the state reading $\sigma_{i}$ in the automaton of $v$. Each node of this trie corresponding to a start of a prefix in $X_{y}$ is added to the set of initial states (see Fig. 2, right). It is easy to see that the constructed NBA is consistent with the sample. To see that it is a SUBA, consider a word $w\in\Sigma^{\omega}$ and assume $w=xy^{\omega}$ for $y$ a shortest period of $w$ and $x$ the shortest prefix reaching such $y$. Note that there is only one such representation of $w$. By the construction of the SUBA, $w$ can only be accepted via a cycle reading $y$ and there is only one such cycle, moreover the cycle can be entered only at the position after reading $x$, and by the suffix-sharing trie there is only one state from which reading $x$ gets to this position. ∎
@@ 5.4. Proving efficient teachability of the informative classes — overview
Because the acceptance conditions differ, $T_{Acc}$ is different for each type of acceptor we consider. In Section 7 we describe the construction of $T_{Acc}$ for acceptors of types IMA and IBA and learning algorithms for acceptors of these types, showing that $\mathbb{IMA}$, $\mathbb{IBA}$, and $\mathbb{ICA}$ are identifiable in the limit using polynomial time and data. In Section 8 we describe the construction of $T_{Acc}$ for acceptors of type IPA and a learning algorithm for acceptors of this type, showing that $\mathbb{IPA}$ is identifiable in the limit using polynomial time and data. In Section 9 we describe the construction of $T_{Acc}$ for acceptors of type IRA and a learning algorithm for acceptors of this type, showing that $\mathbb{IRA}$ and $\mathbb{ISA}$ are identifiable in the limit using polynomial time and data.
@@ 6. The sample $T_{Aut}$ for the automaton
In this section we describe construction of the $T_{Aut}$ part of the sample. We first show that if two states of the automaton are distinguishable, they are distinguishable by words of length polynomial in the number of states of the automaton.
@@ 7. The samples $T_{Acc}$ and learning algorithms for IMA and IBA
The straightforward cases of Muller, Büchi, and coBüchi acceptance conditions are covered in this section. Subsequent sections cover the cases of parity, Rabin, and Street acceptance conditions, which are somewhat more involved.
@@ 7.1. Muller acceptors
If the sample $T$ is consistent with $\mathcal{A}$ and subsumes $T_{L}$, then by Thm. 6.2 the automaton $\mathcal{M}$ returned by $\mathbf{L}_{Aut}$ is isomorphic to the automaton of $\mathcal{A}$. Then by Thm. 7.1 the acceptor returned by $\mathbf{L}_{Acc}^{{\mbox{\scriptsize{IMA}}}}$ with inputs $\mathcal{M}$ and $T$ is an IMA equivalent to $\mathcal{A}$. ∎
@@ 7.2. Büchi  acceptors
If the sample $T$ is consistent with $\mathcal{A}$ and subsumes $T_{L}$ then $\mathbf{L}_{Aut}$ returns an automaton $\mathcal{M}$ isomorphic to the automaton of $\mathcal{A}$ by Thm. 6.2. Then the acceptor returned by $\mathbf{L}_{Acc}^{\mbox{\scriptsize{IBA}}}$ on inputs $\mathcal{M}$ and $T$ is an IBA equivalent to $\mathcal{A}$ by Thm. 7.3. ∎
8. The sample $T_{Acc}$ and learning for IPA
@@ 8.1. A decreasing forest of SCCs of an automaton
After the merge, the roots of $\mathcal{F}$ remain the elements of $\textit{maxSCCs}(S)$. Let $D_{1},\ldots,D_{k}$ be the children of $C$ in $F$, where $D=D_{k}$, and let $E_{1},\ldots,E_{\ell}$ be the children of $D$ in $F$. Because the union of $E_{1},\ldots,E_{\ell}$ is a proper subset of $D=D_{k}$ and the union of $D_{1},\ldots,D_{k}$ is a proper subset of $C$, the union of $D_{1},\ldots,D_{k-1}$ with the union of $E_{1},\ldots,E_{\ell}$ is a proper subset of $C$, therefore the union of the children of $C$ in $\mathcal{F}$ is a proper subset of $C$. Also, the children of $C$ in $\mathcal{F}$ are the maximum SCCs of $(C\setminus\Delta_{\mathcal{F}^{\prime}}(C))$, and no other nodes are affected, so $\mathcal{F}$ is a decreasing forest of SCCs of $\mathcal{M}$ rooted in $S$. ∎
@@ 8.2. Constructing the Canonical Forest and Coloring of a DPA
We describe an algorithm to construct the minStates-forest of $\mathcal{P}$. The roots of the minStates-forest are the elements of $\textit{maxSCCs}(Q)$, each marked as unprocessed. If $C$ is unprocessed, then ${\mathcal{D}}=\textit{maxSCCs}(C\setminus\textit{minStates}(C))$ is computed. If ${\mathcal{D}}$ is empty, $C$ becomes a leaf in the forest, and is marked as processed. Otherwise, $C$ is marked as processed and the elements of ${\mathcal{D}}$ are made the children of $C$ and are marked as unprocessed.
@@
Let $w$ be an $\omega$-word and let $D={\textsl{inf}}(w)$. This is an SCC of the (common) automaton of $\mathcal{P}$ and $\mathcal{P}^{*}$. Let $C=\mathbf{Node}(D,\mathcal{F}^{*}(\mathcal{P}))$. Then $D\cap\Delta(C)\neq\emptyset$, and $\kappa^{*}(D)=\kappa^{*}(C)$, by the definition of $\kappa^{*}$, The $\kappa^{*}$ parity of $C$ is the same as the $\kappa$-parity of $C$, by property (2) of Thm. 8.5. The $\kappa$-parity of $C$ is the same as the $\kappa$-parity of $D$, by property (1) of Thm. 8.5. Thus, the $\kappa^{*}$-parity of $D$ is the same as the $\kappa$-parity of $D$, and $w\in{\llbracket}\mathcal{P}{\rrbracket}$ iff $w\in{\llbracket}\mathcal{P}^{*}{\rrbracket}$. ∎

### v4

@@ 3.3. Polynomial time learners and teachers
A class $\mathbb{C}$ has concise characteristic samples if there exist a polynomial $p(n)$, a teacher $\mathbf{T}$, and a learner $\mathbf{L}$ such that for every $\mathcal{C}\in\mathbb{C}$, $\mathbf{T}(\mathcal{C})$ is a characteristic sample for $\mathcal{C}$ and $\mathbf{L}$, and $|\mathbf{T}(\mathcal{C})|\leq p(\textit{size}(\mathcal{C}))$. Recall that the length of a sample is the sum of length of the strings in it. In Corollary 3.6 we considered just the cardinality of $\mathbf{T}(\mathcal{C})$ and not the length of the words in the sample, but in $|\mathbf{T}(\mathcal{C})|$ the length of the words in the sample matters. Note also that the polynomial bound is in terms of the size of the smallest representation of ${\llbracket}\mathcal{C}{\rrbracket}$. If $\mathbb{C}$ has concise characteristic samples then $\mathbb{C}$ is concisely distinguishable, by Lemma 3.1. In the converse direction, we have the following consequence of Corollary 3.6.
@@
Assume the class $\mathbb{C}$ satisfies (C1), (C2) and (C3) and is concisely distinguishable. Then it has concise characteristic samples for a computable teacher $\mathbf{T}$ and a computable learner $\mathbf{L}$.
@@
As has been observed before [Pit89], if (P1) and (P2) are satisfied, this definition is superfluous because the learning algorithm of identification by enumeration used in the proof of Theorem 3.2 can be modified to run in polynomial time and the teacher adjusted appropriately to create a (potentially ridiculously large) characteristic sample. This observation motivates the introduction of the criterion of identification in the limit with polynomial time and data (described in Section 3.4). The argument for this observation follows.
@@ 3.5. Relation to learning with equivalence and membership queries
Suppose the class $\mathbb{C}$ satisfies properties (P1), (P2) and (P3). If $\mathbb{C}$ is polynomially learnable using membership and equivalence queries, then it is also efficiently teachable.
@@ 4.2. Negative results for nondeterministic classes
The classes $\mathbb{NBA}$, $\mathbb{NPA}$, $\mathbb{NMA}$, $\mathbb{NCA}$ do not have concise characteristic sets [AFS20]. The proof is by constructing a family of languages $\{L_{n}\}_{n\in\mathbb{N}}$ with an acceptor of size quadratic in $n$ for which at least one word of length at least exponential in $n$ must be included in any characteristic sample for $L_{n}$. [^7] A negative result regarding query learning of $\mathbb{NBA}$, $\mathbb{NPA}$ and $\mathbb{NMA}$ was obtained by Angluin et al. [AAF20]. That result makes a plausible assumption of cryptographic hardness, which is not required here. Since an NBA (resp. NCA) is a special case of NRA (resp. NSA) the same is true for $\mathbb{NRA}$ and $\mathbb{NSA}$.
@@ 4.3. Consequences of membership and equivalence algorithms
Let $\mathbf{T}$ be a teacher witnessing the efficient teachability of M2MAs. The teacher $\mathbf{T}^{\prime}$ with input a SUBA $\mathcal{C}$ generates the characteristic sample $\mathbf{T}(r(\mathcal{C}))$, to which it adds one example $(x,b)$ such that $x=\mathcal{C}$ and $b=1$ iff $x\in{\llbracket}\mathcal{C}{\rrbracket}$. The learner $\mathbf{L}$ with input $T$ searches for the least $x$ (if any) such that $(x,b)\in T$, $r(x)$ is consistent with $T$, and $\mathbf{T}(r(x))\subseteq T$. If such an $x$ is found, it is output; otherwise, $\mathbf{L}$ constructs and outputs a SUBA consistent with $T$. A procedure constructing a default SUBA acceptor that agrees with a given sample $T$ is given in the proof of Proposition 5.2. ∎
@@ 5.1. Duality
There are reductions of the problem of efficient teachability between $\mathbb{IBA}$ and $\mathbb{ICA}$ and between $\mathbb{IRA}$ and $\mathbb{ISA}$, using the duality between these types of acceptors. Consequently we focus on the classes $\mathbb{IBA}$, $\mathbb{IPA}$, $\mathbb{IRA}$ and $\mathbb{IMA}$ in what follows. [^8] The results regarding the classes $\mathbb{IBA}$ (and $\mathbb{ICA}$), $\mathbb{IPA}$ and $\mathbb{IMA}$ were obtained in [AFS20]; here we extend them to the classes $\mathbb{IRA}$ (and $\mathbb{ISA}$). Results for identifiability in the limit using polynomial time and data (but not efficient teachability) of the classes $\mathbb{IRA}$ (and $\mathbb{ISA}$) have also been provided in [BL21] using a different algorithm.
@@
Let $\mathcal{A}$ be an ICA. Because $\mathcal{A}$ is deterministic and complete, if we let $\mathcal{A}^{\prime}$ denote the IBA with the same components as $\mathcal{A}$, then $\mathcal{A}^{\prime}$ accepts the complement of the language $\mathcal{A}$, by Claim 2.2 (4).
@@
The same conversion may be done with acceptors of types IRA and ISA, by Claim 2.2 (5). ∎
@@ 5.2. The default acceptor
Given a sample $T$, we initialize $U$ to be the empty set, and for every word $w_{i}$ in the sample, we find the shortest prefix $u_{i}$ of $w_{i}$ that distinguishes it from all other examples in $T$ and add it to $U$ (see Figure 2, middle). We arrange the finite words in $U$ in a trie in the usual manner. We add self-loops on each $\sigma\in\Sigma$ to the leaves of the trie (see Figure 2, left). This deterministic automaton is termed the prefix-tree automaton [OG92]. If the automaton is incomplete, we add a new dead state with self-transitions on each $\sigma\in\Sigma$, and define all undefined transitions to go to the dead state. Recall that by Proposition 4.1 the length of a prefix distinguishing two examples $u_{1}(v_{1})^{\omega}$ and $u_{2}(v_{2})^{\omega}$ is polynomially bounded by the length of the examples. It follows that the prefix-tree automaton can be constructed in time polynomial in the length of the sample $T$.
@@
For a DBA, the acceptance condition $F$ consists of all the trie-leaf states that are prefixes of positive examples in $T$ (see Figure 2, left). For a DMA, the acceptance condition consists of $\{\{q\}\mid q\in F\}$. For a DCA, the acceptance condition consists of the dead state (if one was added) and all the trie-leaf states that are not prefixes of positive examples in $T$. The DBA thus constructed may be transformed to a DPA or a DRA using Claim 2.2 (1) or Claim 2.2 (2), respectively, and the DCA may be transformed to a DSA using Claim 2.2 (3).
@@
For a SUBA the construction is different. We first construct a set $V$ of shortest periods of positively labeled examples, by going iteratively over the examples $w_{1},w_{2},\ldots$ and proceeding as follows. For $w_{i}=u(v)^{\omega}$ if $v_{i}$ is a shortest period of $w_{i}$, and none of the rotations of $v_{i}$ is in $V$ we add $v_{i}$ to $V$. Then for every period $v=\sigma_{1}\sigma_{2}\ldots\sigma_{k}$ in $V$ we construct an automaton with $k$ states $s_{0},s_{1},\ldots,s_{k-1}$ and transitions $(s_{i},\sigma_{i+1},s_{i^{\prime}})$ for $0\leq i<k$ where $i^{\prime}=(i+1)\mod k$. We designate $s_{0}$ as an accepting state. Finally, for each positively labeled word $w$ we look for the shortest prefix $x$ of $w$ such that $w=x(y)^{\omega}$ for some rotation $y=\sigma_{i}\ldots\sigma_{k}\sigma_{1}\ldots\sigma_{i-1}$ of a period $v=\sigma_{1}\sigma_{2}\ldots\sigma_{k}\in V$. For a rotation $y$ of $v\in V$, let $X_{y}$ be the set of such shortest prefixes. The prefixes in $X_{y}$ are arranged in a suffix-sharing trie and the trie is connected to the automata constructed for the periods by landing in the state reading $\sigma_{i}$ in the automaton of $v$. Each node of this trie corresponding to a start of a prefix in $X_{y}$ is added to the set of initial states (see Figure 2, right). It is easy to see that the constructed NBA is consistent with the sample. To see that it is a SUBA, consider a word $w\in\Sigma^{\omega}$ and assume $w=xy^{\omega}$ for $y$ a shortest period of $w$ and $x$ the shortest prefix reaching such $y$. Note that there is only one such representation of $w$. By the construction of the SUBA, $w$ can only be accepted via a cycle reading $y$ and there is only one such cycle, moreover the cycle can be entered only at the position after reading $x$, and by the suffix-sharing trie there is only one state from which reading $x$ gets to this position. ∎
@@ 5.4. Proving efficient teachability of the informative classes — overview
Because the acceptance conditions differ, $T_{Acc}$ is different for each type of acceptor we consider. In Section 7 we describe the construction of $T_{Acc}$ for acceptors of types IMA and IBA and learning algorithms for acceptors of these types, showing that $\mathbb{IMA}$, $\mathbb{IBA}$ and $\mathbb{ICA}$ are identifiable in the limit using polynomial time and data. In Section 8 we describe the construction of $T_{Acc}$ for acceptors of type IPA and a learning algorithm for acceptors of this type, showing that $\mathbb{IPA}$ is identifiable in the limit using polynomial time and data. In Section 9 we describe the construction of $T_{Acc}$ for acceptors of type IRA and a learning algorithm for acceptors of this type, showing that $\mathbb{IRA}$ and $\mathbb{ISA}$ are identifiable in the limit using polynomial time and data.
@@ 6. The sample $T_{Aut}$ for the automaton
In this section we describe the construction of the $T_{Aut}$ part of the sample. We first show that if two states of the automaton are distinguishable, they are distinguishable by words of length polynomial in the number of states of the automaton.
@@ 7. The samples $T_{Acc}$ and learning algorithms for IMA and IBA
The straightforward cases of Muller, Büchi and coBüchi acceptance conditions are covered in this section. Subsequent sections cover the cases of parity, Rabin and Street acceptance conditions, which are somewhat more involved.
@@ 7.1. Muller acceptors
If the sample $T$ is consistent with $\mathcal{A}$ and subsumes $T_{L}$, then by Theorem 6.2 the automaton $\mathcal{M}$ returned by $\mathbf{L}_{Aut}$ is isomorphic to the automaton of $\mathcal{A}$. Then by Theorem 7.1 the acceptor returned by $\mathbf{L}_{Acc}^{{\mbox{\scriptsize{IMA}}}}$ with inputs $\mathcal{M}$ and $T$ is an IMA equivalent to $\mathcal{A}$. ∎
@@ 7.2. Büchi  acceptors
If the sample $T$ is consistent with $\mathcal{A}$ and subsumes $T_{L}$ then $\mathbf{L}_{Aut}$ returns an automaton $\mathcal{M}$ isomorphic to the automaton of $\mathcal{A}$ by Theorem 6.2. Then the acceptor returned by $\mathbf{L}_{Acc}^{\mbox{\scriptsize{IBA}}}$ on inputs $\mathcal{M}$ and $T$ is an IBA equivalent to $\mathcal{A}$ by Theorem 7.3. ∎
8. The sample $T_{Acc}$ and the learning algorithm for IPA
@@ 8.1. A decreasing forest of SCCs of an automaton
After the merge, the roots of $\mathcal{F}$ remain the elements of $\textit{maxSCCs}(S)$. Let $D_{1},\ldots,D_{k}$ be the children of $C$ in $F$, where $D=D_{k}$, and let $E_{1},\ldots,E_{\ell}$ be the children of $D$ in $F$. Because the union of $E_{1},\ldots,E_{\ell}$ is a proper subset of $D=D_{k}$ and the union of $D_{1},\ldots,D_{k}$ is a proper subset of $C$, the union of $D_{1},\ldots,D_{k-1}$ with the union of $E_{1},\ldots,E_{\ell}$ is a proper subset of $C$, therefore the union of the children of $C$ in $\mathcal{F}$ is a proper subset of $C$. Also, the children of $C$ in $\mathcal{F}$ are the maximum SCCs of $C\setminus\Delta_{\mathcal{F}^{\prime}}(C)$, and no other nodes are affected, so $\mathcal{F}$ is a decreasing forest of SCCs of $\mathcal{M}$ rooted in $S$. ∎
@@ 8.2. Constructing the Canonical Forest and Coloring of a DPA
We describe an algorithm to construct the minStates-forest of $\mathcal{P}$. The roots of the minStates-forest are the elements of $\textit{maxSCCs}(Q)$, each marked as unprocessed. The following procedure is repeated until all elements are marked as processed. If $C$ is unprocessed, then ${\mathcal{D}}=\textit{maxSCCs}(C\setminus\textit{minStates}(C))$ is computed. If ${\mathcal{D}}$ is empty, $C$ becomes a leaf in the forest, and is marked as processed. Otherwise, $C$ is marked as processed and the elements of ${\mathcal{D}}$ are made the children of $C$ and are marked as unprocessed.
@@
Let $w$ be an $\omega$-word and let $D={\textsl{inf}}(w)$. This is an SCC of the (common) automaton of $\mathcal{P}$ and $\mathcal{P}^{*}$. Let $C=\mathbf{Node}(D,\mathcal{F}^{*}(\mathcal{P}))$. Then $D\cap\Delta(C)\neq\emptyset$, and $\kappa^{*}(D)=\kappa^{*}(C)$, by the definition of $\kappa^{*}$, The $\kappa^{*}$ parity of $C$ is the same as the $\kappa$-parity of $C$, by property (2) of Theorem 8.5. The $\kappa$-parity of $C$ is the same as the $\kappa$-parity of $D$, by property (1) of Theorem 8.5. Thus, the $\kappa^{*}$-parity of $D$ is the same as the $\kappa$-parity of $D$, and $w\in{\llbracket}\mathcal{P}{\rrbracket}$ iff $w\in{\llbracket}\mathcal{P}^{*}{\rrbracket}$. ∎